OpenStudy (anonymous):

i need help solving 3x/5=2x/5+10

OpenStudy (anonymous):

what is the LCD?

OpenStudy (anonymous):

um what is that ???

OpenStudy (anonymous):

3x/5=2x/5+10 3x=2x/5+10 3x=2x/5*5+10*5 3x=2x*1+10*5 3x=2x+10*5 3x-2x=50 x=50 (I think??)

OpenStudy (anonymous):

Katherine i had got the same thing when i wrote it out but i just wasnt sur thanks

OpenStudy (anonymous):

No problemo! :) Glad I could be of assistance!